OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side of the triangle. Round to the nearest tenth, if necessary. a = ?, b = 5, c = 8 6.3 6.2 6.0 The sides of the second triangle are: 24,7,c 20 25 26 30

Parth (parthkohli):

a^2 + b^2 = 8^2 a^2 + 25 = 64 a^2 = 49 a = 7 Just like that the answer of 2nd is 25